OpenStudy (anonymous):

integrate tan to the power 5X

OpenStudy (anonymous):

\[\tan(x)^{5x}\] Is that the expression?

OpenStudy (anonymous):

variable in the exponent, consider logarthmic manipulation: integrate 5x ln tan theta

OpenStudy (anonymous):

Yeah, you didn't say tan of what

OpenStudy (anonymous):

I'd guess (in the absence of any variable except x) that it is: \[\tan^{5}x \] But of course, I could be wrong.

OpenStudy (anonymous):

Where is oneprince?

OpenStudy (anonymous):

The prince has stepped in his concubine to satisfy one of his maidens

OpenStudy (anonymous):

INweton integrate it.that is the correct question

OpenStudy (anonymous):

Re-write it as \[\tan x(\sec^2x -1)(\sec^2x -1) \] Expand and do it term-by-term. Reduction formulae seems unnecessary, and I can;t see anything quicker right now :@.

OpenStudy (anonymous):

You simply let tan x be u and integrate u^5

OpenStudy (anonymous):

INewton continue

OpenStudy (anonymous):

OK, I'll do a little more. That expand to \[\tan x \sec^4x - 2\tan x\ \sec^2x + \tan x \] And note that \[\frac{\mathbb{d}}{\mathbb{d}x}\frac{1}{n}\ \sec^nx =\tan x\ \sec^nx \] Again, probably something far quicker.

OpenStudy (anonymous):

Chaguanas, that doesn't work well because you'll end up with a factor \(cos^2x\) from the substitution of dx.

OpenStudy (anonymous):

I think it does work (sec^2x = 1 + tan^2x) but it's not MUCH nicer. Again, pretty tired so may be wrong.

OpenStudy (anonymous):

Actually, looking at what comes out I'd wager not nicer full stop.

OpenStudy (anonymous):

Alternative method: tables: integration tan^u du=(1/(n-1)) tan^(n-1) u - integral tan^(n-2) u du